# An Efficient Learning Control Framework With Sim-to-Real for String-Type Artificial Muscle-Driven Robotic Systems

### v2

@@ Abstract
Machine learning for robot control, artificial muscle, deep reinforcement learning, sim-to-real.
@@ I Introduction
Artificial muscles, designed to mimic biological counterparts, generate mechanical forces and movements in response to external stimuli like electrical or chemical signals. Well known for their high power-to-weight ratios and inherent compliance [1], artificial muscles offer a unique combination of strength and flexibility conducive to versatile and adaptive robot designs. Over the past decade, robotic systems driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including biomimetic robots [2, 3, 4], robot manipulators and grippers [5, 6], soft robots [7, 8], and many others [9, 10, 11].
Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains a challenge due to their highly nonlinear and complex dynamics. While previous research efforts have primarily centered around structural design and open-loop motion performance evaluation [12, 13, 14], recent studies have begun to investigate closed-loop control strategies with a majority exploring proportional-integral-derivative (PID) controllers [15, 16, 17]. However, PID controllers often fall short in consistently delivering desired control performance, especially in complex robotic systems.
With the rapid development of machine learning techniques, learning-based controllers have emerged as an increasingly popular avenue in the muscle-driven robotics community. For instance, N. Nikdel et al. [6] developed a neural network-based model predictive controller for robotic manipulators actuated by shape memory alloys (SMAs). M. Giorelli et al. [18] established the inverse kinematics of cable-driven soft robotic arms using feedforward neural networks. T. Yang et al. [19] proposed a reinforcement learning control strategy for a cuttlefish-like soft robot driven by dielectric elastomer (DE) membranes. Among these approaches, deep reinforcement learning (DRL) stands out for its remarkable performance across diverse and complex tasks, offering a universal control solution. An increasing number of researchers have leveraged DRL algorithms to address control challenges in artificial muscle-driven robots, spanning applications ranging from dexterous grasping of robotic hands[20] to agile motion control of soft robotic manipulators [21, 22], and various others [23, 24, 25, 26].
In most of existing studies [20, 22, 24, 23, 26], the standard procedure of applying DRL control in the targeted robotic systems typically follows the schematic illustrated in Fig. 1. Initially, it is crucial to identify the relevant parameters of the artificial muscles and formulate the system dynamics model to establish a simulation environment. Subsequently, the DRL algorithms undergo training within simulation to develop a control policy. Finally, the acquired control policy is deployed on the robotic platform to accomplish various control objectives. Despite these advancements, challenges persist in the application of DRL control in the targeted robotic systems.
Figure 1: Schematic of standard procedure of applying DRL control in artificial muscle-driven robotic systems. Three main steps include 1) dynamic modeling with system identification to establish simulation environment; 2) training DRL agent within the environment to develop a control policy; and 3) deploying the trained policy onto real-world robotic systems for testing.
I-1 Muscle Sim-to-real Gap
For the robotic systems driven by artificial muscles, modeling errors primarily come from the following three aspects. First, the dynamic models of artificial muscles themselves are approximate and inaccurate [27, 28]. Second, inevitable errors occur in the system identification process of the muscles. Third, different muscle strings typically have different model parameters [29, 30], and conducting system identification for all muscle strings is time-consuming and labor-intensive. These factors jointly contribute to inaccurate modeling of artificial muscles, resulting in a significant sim-to-real gap for the entire robotic system. Consequently, control strategies that perform well in simulation may not seamlessly transfer to real-world applications [21, 22]. Despite being labor-intensive, some researchers [19, 25] have attempted to directly train DRL algorithms on real robots. However, given the fragile nature of string-type artificial muscles [1], this approach lacks universality and safety guarantees. Hence, addressing the muscle sim-to-real gap for DRL-based controllers is imperative.
I-2 Low Learning Efficiency
In reinforcement learning, an agent continually interacts with the environment to acquire a good policy through learning from the interaction data [31]. Due to the complexity and high-dimensional nature of deep neural networks, DRL algorithms usually require a large amount of training data and iterations for convergence towards the desired policy. While training in simulation remains feasible, many studies have noted that the low sample efficiency often leads to significantly prolonged training times [23, 26, 32]. Moreover, the complex dynamic models of muscle-driven robots exacerbate this issue. Thus, enhancing training data efficiency and accelerating convergence speed are of great importance.
We observe a scarcity of research addressing the aforementioned issues within the domain of artificial muscle-driven robots. Our prior research [24] aimed to tackle the challenge of low learning efficiency. However, its experimental findings were solely derived from simulations, without addressing the crucial muscle sim-to-real gap. This paper presents a DRL-based control strategy with improved data efficiency and transferability to real-world applications. It is worth noting that in recent years, there have been some efforts to bridge the sim-to-real gap of artificial muscles. Many researchers [27, 28, 33, 34, 35] have proposed more accurate yet significantly more complex dynamic models. In this paper, however, we seek to directly improve the performance of DRL-based controllers by employing a straightforward sim-to-real transfer technique, thereby bypassing the complicated process of mathematical modeling and parameter identification for artificial muscle strings. We evaluate our learning control frame work on two robotic systems powered by artificial muscle strings, including a biomimetic robotic eye driven by super-coiled polymers (SCPs) [36, 29], and a parallel robotic wrist actuated by twisted-coiled actuators (TCAs) [30, 37].
@@ I-2 Low Learning Efficiency
The main contributions of this paper are as follows. To the best of our knowledge, this paper proposes the first efficient learning control framework designed for string-type artificial muscle-driven robotic systems. Specifically, we introduce a novel augmentation method for state-based data, which transforms a single transition data into several. We leverage an easy-to-implement PID controller to generate trajectories for early-stage training of DRL agents. The training results in simulation show that our proposed methods improve the data efficiency of the DRL algorithm in targeted robotic systems. Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics. The experimental results demonstrate a great improvement in real-world control performance through the application of this method in muscle-driven robotic systems. We believe that the proposed learning control framework can be utilized to address a variety of control tasks across similar robotic systems.
@@
The remainder of this paper is organized as follows. Section II presents basic information about leveraging reinforcement learning within the robotic systems of interest. Section III offers a concise overview of the robotic platforms utilized in this study. Section IV presents our proposed efficient learning control framework in detail. Subsequently, Section V presents the experimental results of the proposed method, encompassing training outcomes in simulation and control performances in real-world scenarios. Finally, concluding remarks and current limitations are provided in Section VI.
II Preliminaries
@@
In this section, we model the control problem of targeted robotic systems as a reinforcement learning problem by Markov Decision Process (MDP), which is a mathematical framework that describes the interaction between the DRL agent and the environment [31]. The MDP in artificial muscle-driven robots consists of the following tuple, i.e.,
$(\mathcal{S},\mathcal{A},\mathcal{R},f,\gamma).$ (1)
@@
  • $\mathcal{S}$ is the state space. At each time step $t$, the state vector $s_{t}\in\mathcal{S}$ generally includes the current state variables of the robotic system and the target positions and/or orientations. In this paper, the motion state variables of the robotic systems are collectively defined as vector $\bm{x}$, which contains the Euler angles, angular velocities, and other relevant motion states. The system output vector, containing the variables to be controlled, is denoted by $\bm{y}$. Correspondingly, the target system output is denoted by $\bm{y}^{*}$. In this way, the state vector $s_{t}$ is represented as
    $s_{t}=(\bm{x}_{t},\bm{y}^{*}).$ (2)
@@
  • $\mathcal{A}$ is the action space, where each action $a_{t}\in\mathcal{A}$ is generated by the DRL agent based on the current state $s_{t}$ and its own policy. The action vector typically consists of the actuating voltage applied to each muscle or the voltagel difference between a pair of antagonistic muscles.
  • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}$, i.e.,
    $r_{t}=\mathcal{R}(s_{t},a_{t}).$ (3)
@@
    The design of reward function plays a vital role in the learning process of the DRL agents, which learn the optimal policy under the guidance thereby. In this paper, we develop a generalized quadratic-form reward function design for the robotic systems of interest:
    $r_{t}=-(e^{T}_{t}Q_{e}e_{t}+a^{T}_{t}R_{a}a_{t})+r_{\text{bonus}}$ (4)
@@
    where $e_{t}=|\bm{y}^{*}-\bm{y}_{t}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively. A bonus reward $r_{\text{bonus}}$ is given when the system tracking error achieves a sufficiently small threshold. We add this bonus reward because, for the robotic systems discussed in this paper, maintaining at the target state is generally more challenging than reaching it.
  • $f$ represents the state transition function, which describes the likelihood of a transition from state $s$ to $s^{\prime}$ when an action $a$ is taken.
  • $\gamma\in[0,1]$ is the discount factor, which serves to balance the influence of future rewards on the current state. In all experiments conducted in this paper, we set $\gamma$ to a constant value of $\gamma=0.99$.
III Experimental Platforms
@@
This section gives a concise overview of the testing platforms utilized in this paper, including an SCP-driven robotic eye (see Fig. 2) and a TCA-driven parallel robotic wrist (see Fig. 3). We briefly describe the hardware design, state space, action space, and reward function for each system.
@@
III-A Robotic Eye
The high occurrence of unsuccessful corrective surgeries for ocular motor disorders underscores the need for a muscle-driven robotic eye platform to assist ophthalmologists in gaining deeper insights into the biomechanics of human eye movement. In our previous work [36], we proposed a prototype of a 2-DOF robotic eye, and derived the geometric and dynamic models specific to this prototype. As illustrated in Fig. 2, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC). The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles by regulating the SCP muscles, namely foveation control in the eye movement study.
The eyeball’s center of rotation, denoted as $\bm{O}\in\mathbb{R}^{3}$, remains fixed at the origin within the reference frame $O_{xyz}$. Here, the axes $x$, $y$, and $z$ correspond to the transverse, sagittal, and vertical axes of the eyeball, respectively. Rotations around these axes are expressed as pitch $\theta$, roll $\phi$, and yaw $\psi$. The robotic eye system’s state variables are represented as
$\bm{x}=[\theta,\dot{\theta},\psi,\dot{\psi},\Delta T_{1},\Delta T_{2},\Delta T_{3},\Delta T_{4}]$ (5)
where $\theta,\dot{\theta},\psi,\dot{\psi},$ represent the pitch, yaw angles, and their respective angular velocities for control. Additionally, $\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{0}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
Figure 3: The parallel robotic wrist prototype actuated by three TCAs.
Focusing primarily on the robotic eye’s orientation, the system’s output vector is denoted as $\bm{y}=[\psi,\dot{\psi},\theta,\dot{\theta}]$. Due to the difficulty and poor accuracy of measuring the temperature of each muscle in real time, the state vector of the DRL agent is defined by
$s_{t}=[\theta_{t},\dot{\theta}_{t},\psi_{t},\dot{\psi}_{t},\theta^{*},\psi^{*}].$ (6)
@@ III-A Robotic Eye
$a_{t}=[$ $-V_{1}+V_{2},-V_{3}+V_{4}]$ (7)
s.t. $V_{1}\cdot V_{2}=V_{3}\cdot V_{4}=0$
$0\leq V_{i}\leq 10$
The constraint requires that only one muscle is activated at each time step in the pairs $(m_{1},m_{2})$ and $(m_{3},m_{4})$. The reward function follows the structure of (4), where the output and control weight matrices are set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. The bonus reward is granted when the tracking error is sufficiently small, given by

### v3

@@ Abstract
Reinforcement learning for robot control, artificial muscle, sim-to-real, sample efficiency.
@@ I Introduction
Artificial muscles, designed to mimic biological counterparts, generate mechanical forces and movements in response to external stimuli like electrical or chemical signals. Well known for their high power-to-weight ratios and inherent compliance[1], artificial muscles offer a unique combination of strength and flexibility conducive to versatile and adaptive robot designs. Over the past decade, robots driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including but not limited to biomimetic robots[2, 3], robot manipulators[4, 5], soft robots[6, 7], medical robot[8], origami robots[9], and marine robots[10].
Previous research efforts have primarily centered around structural design and open-loop motion performance evaluation[2, 3, 4, 5, 6, 7, 8, 9, 10]. Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains challenging due to their highly nonlinear and complex dynamics. Traditional control strategies, such as proportional-integral-derivative (PID) control[11, 12, 13], energy-based nonlinear control[14], and sliding mode control[15, 16] have been investigated to mitigate these challenges. Although these approaches have demonstrated success in specific tasks, a comprehensive, unified control strategy for string-type artificial muscle-driven robots is still lacking. As deep reinforcement learning (DRL) rapidly develops, learning-based controllers[17, 18, 19, 20, 21] have emerged as a promising alternative, offering the potential to overcome the limitations of conventional methods. The standard procedure of applying DRL in the targeted robotic systems typically follows the schematic illustrated in Fig. 1. Initially, identifying the relevant parameters and establishing a simulation environment enable a safe and efficient training process. The DRL algorithms are then trained within this simulation environment to develop an effective control policy. Finally, the learned control policy is transferred to the physical robotic platform to achieve the desired control objectives.
Figure 1: Schematic of standard procedure of applying DRL control in artificial muscle-driven robotic systems.
However, several challenges persist in the applications of DRL to string-type artificial muscle-driven robots. One major issue is the sim-to-real gap, which arises from inaccuracies in dynamic models, system identification errors, and variability in muscle string parameters[22, 23]. These factors often result in control strategies that demonstrate strong performance in simulation but struggle to transfer effectively to real-world applications. Another challenge is the low learning efficiency of DRL algorithms, which require substantial amounts of training data and iterations due to the complexity of neural networks and the trial and error learning process[24]. Addressing these two challenges is crucial for advancing control technologies in artificial muscle-driven robots, which are increasingly employed in soft robotics and bio-inspired systems.
@@
This paper presents a DRL-based control strategy designed to enhance the data efficiency and real-world applicability of controllers for artificial muscle-driven robots. The proposed framework is evaluated on two robotic systems actuated by artificial muscle strings: a biomimetic robotic eye driven by super-coiled polymers (SCPs)[25, 26] and a parallel robotic wrist actuated by twisted-coiled actuators (TCAs)[27, 28]. Specifically, we introduce a novel augmentation method for state-based data, which transforms a single transition data tuple into several, and we leverage an easy-to-implement PID controller to generate trajectories for early-stage training of DRL agents. Compared to the baseline algorithm (i.e., Soft Actor Critic [29]), the required training data is reduced by approximately $60\%$ in the targeted robotic systems. Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics. The experimental results demonstrate a significant improvement in real-world control performance with the averaged error reduced by $36.3\%$ through the application of this method.
@@
The remainder of this paper is organized as follows. Section II provides a review of the relevant literature. Section III introduces the basic concepts of applying reinforcement learning to the robotic systems studied. Section IV offers a concise overview of the robotic platforms utilized in this study. In Section V, we present a detailed description of our proposed efficient learning control framework. Section VI and Section VII discuss the experimental results, including training outcomes in simulation and control performances in real-world scenarios, respectively. Finally, concluding remarks and current limitations are discussed in Section VIII.
II Related Work
The control of artificial muscle-driven robotic systems using learning-based methods has garnered significant attention in recent years[30, 31, 20, 19, 32, 21, 18, 33]. Among these approaches, DRL has demonstrated effectiveness in handling diverse and complex tasks, positioning it as a promising control solution for a wide range of applications. An increasing number of researchers have leveraged DRL algorithms to address control challenges in artificial muscle-driven robots, spanning applications ranging from dexterous grasping of robotic hands[17] to agile motion control of soft robotic manipulators[33, 18], and other related applications[32, 21]. Many of these studies have noted that the low sample efficiency often leads to prolonged training time[20, 21], but few have focused specifically on improving training efficiency. Several existing state-of-the-art (SOTA) reinforcement learning (RL) methods may be applicable[34, 35, 36, 37], but they have not been evaluated within the context of the robotic systems we are addressing. The unique nonlinear and slow response characteristics of artificial muscle may present challenges that are not adequately addressed by these existing methods. Our previous work[19] explored low learning efficiency problem but was limited to simulation-based experiments.
Due to the sim-to-real gap, control strategies that perform well in simulation may not seamlessly transfer to real-world applications[33, 18]. Despite being labor-intensive, some researchers[31, 32] have attempted to directly train DRL algorithms on real robots. However, the artificial muscles considered in this paper are all thermally actuated, which inherently leads to energy inefficiency and limited cycling rates[38]. Continuous rapid transitions between heating and cooling often cause irreversible damage to the muscles, making direct training on real robots unsafe. It is worth noting that in recent years, there have been some efforts to bridge the sim-to-real gap of artificial muscles. Many researchers[22, 23, 39] have proposed more accurate yet significantly more complex dynamic models. In this paper, however, we aim to directly improve the performance of DRL-based controllers by employing a sim-to-real transfer technique, thereby reducing the need for complicated mathematical modeling and precise parameter identification for artificial muscles.
III Preliminaries
@@
In this section, we model the control problem of targeted robotic systems as a reinforcement learning problem by Markov Decision Process (MDP), which consists of the tuple $(\mathcal{S},\mathcal{A},\mathcal{R},f,\gamma)$.
@@
  • $\mathcal{S}$ is the state space. The state vector $s_{t}\in\mathcal{S}$ generally includes the current state variables of the robotic system and the target positions and/or orientations. In this paper, the motion state variables of the robotic systems are collectively defined as vector $\bm{x}$. The system output vector, containing the variables to be controlled, is denoted by $\bm{y}$. Correspondingly, the target system output is denoted by $\bm{y}^{*}$. In this way, the state vector $s_{t}$ is represented as $s_{t}=(\bm{x}_{t},\bm{y}^{*})$.
@@
  • $\mathcal{A}$ is the action space, where each action $a_{t}\in\mathcal{A}$ is generated by the DRL agent based on the current state $s_{t}$ and its own policy.
@@
  • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}=\mathcal{R}(s_{t},a_{t})$. The design of reward function plays a vital role in the learning process of the DRL agents, which learn the optimal policy under the guidance thereby. In this paper, we develop a generalized quadratic-form reward function designed for the robotic systems of interest:
    $r_{t}=-(e^{\top}_{t}Q_{e}e_{t}+a^{\top}_{t}R_{a}a_{t})+r_{\text{bonus}},$ (1)
@@
    where $e_{t}=|\bm{y}^{*}-\bm{y}_{t}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively. A bonus reward $r_{\text{bonus}}$ is given when the system tracking error falls below a sufficiently small threshold.
  • $f$ represents the state transition function and $\gamma$ is the discount factor, which is set to $0.99$ in our experiments.
Figure 2: Illustrations of SCP and TCA muscles. (a) Crimp fixtures for the SCP muscle ends (left), and a close-up shot of the fabricated SCP muscle (right); and (b) conductive ring with SMA wire wrapped around for the TCA muscle ends (left), and a close-up shot of the fabricated TCA muscle (right).
IV Experimental Platforms
@@
This section presents an overview of the actuators and robotic platforms used in this paper, including SCP and TCA artificial muscles (Fig. 2), a 2-DOF robotic eye (Fig. 3) and a parallel robotic wrist (Fig. 4). For each system, we briefly describe the hardware design, state space, action space, and reward function.
IV-A Artificial Muscles
Numerous fabrication methods for SCP and TCA muscles have been investigated in the literature[25, 26, 27, 28]. Specifically, for our robotic eye design, the SCP actuator is fabricated by sequentially twisting, coiling, and double-backing using a 145 cm long silver-plated conductive nylon thread (Statex’s Shieldex 117/17 dtex 2 ply HC+B). As illustrated in Fig. 2(a), the ends of the SCP muscles are crimped using fold-over silver/copper bead crimps. They are connected to N-Channel enhancement MOSFETs (FQP30N06L) which provide electric power to the muscles. Application of voltage induces Joule heating on the actuator which causes the coils to expand laterally, thus generating contracting force.
Regarding our TCA muscles, the Shape Memory Alloy (SMA) wires (0.18 mm diameter) serve as the conductive threads and skeletons. Highly elastic spandex fibers (creora spandex, power fit, 840 denier) are tightly wound around the SMAs, and the resulting composite fibers are then twisted into a coil configuration. As shown in Fig. 2(b), both ends of the SMA wire are wrapped around conductive rings, which are also connected to MOSFETs. When voltage potential is applied to the rings, the skeleton generates Joule heat and the surrounding spandex fibers will contract in longitudinal direction, which leads to the untwisting of the compound twisted fiber structure. The TCA’s coiled configuration converts this untwisting motion into linear contraction. Additionally, the SMA wire not only provides Joule heating, but also contributes additional contraction force owing to its shape memory effect.
@@ IV-A Artificial Muscles
IV-B Robotic Eye
The high occurrence of unsuccessful corrective surgeries for ocular motor disorders underscores the need for a muscle-driven robotic eye platform to assist ophthalmologists in gaining deeper insights into the biomechanics of human eye movement. In our previous work[25], we proposed a prototype of a 2-DOF robotic eye and derived the geometric and dynamic models specific to this prototype. As illustrated in Fig. 3, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC). The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles by regulating the SCP muscles, namely foveation control in the eye movement study.
The eyeball’s center of rotation, denoted as $\bm{O}\in\mathbb{R}^{3}$, remains fixed at the origin within the reference frame $O_{\mathrm{xyz}}$. Here, the axes $\mathrm{x}$, $\mathrm{y}$, and $\mathrm{z}$ correspond to the transverse, sagittal, and vertical axes of the eyeball, respectively. Rotations around these axes are expressed as pitch $\theta$, roll $\phi$, and yaw $\psi$. The robotic eye system’s state variables are represented as
$\bm{x}=[\theta,\dot{\theta},\psi,\dot{\psi},\Delta T_{1},\Delta T_{2},\Delta T_{3},\Delta T_{4}]^{\top},$ (2)
where $\theta,\dot{\theta},\psi,\dot{\psi},$ represent the pitch, yaw angles, and their respective angular velocities for control. $\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{0}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
Figure 4: The parallel robotic wrist prototype actuated by three TCA muscles.
Focusing primarily on the robotic eye’s orientation, the system’s output vector is denoted as $\bm{y}=[\psi,\dot{\psi},\theta,\dot{\theta}]^{\top}$. Due to the difficulty and poor accuracy of measuring the temperature of each muscle in real time, the state vector of the DRL agent is defined by
$s_{t}=[\theta_{t},\dot{\theta}_{t},\psi_{t},\dot{\psi}_{t},\theta^{*},\psi^{*}]^{\top}.$ (3)
@@ IV-B Robotic Eye
$a_{t}=[$ $-V_{1}+V_{2},-V_{3}+V_{4}]^{\top}$ (4)
[…]
The constraint requires that only one muscle is activated at each time step in the pairs $(m_{1},m_{2})$ and $(m_{3},m_{4})$. The reward function follows the structure of (1), where the output and control weight matrices were set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. They were chosen based on empirical tuning to balance the trade-off between output error and control effort. The bonus reward is granted when the tracking error is sufficiently small, given by
@@
IV-C Robotic Wrist
Research into artificial muscle-driven robotic wrist joints is of great importance to several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields. We designed a parallel robotic wrist prototype driven by three equally spaced compound TCAs[27] as illustrated in Fig. 4. The base plate remains stationary, while the upper plate, also known as the end effector, moves via three muscle strings. Linkages between the two plates are constructed to support a two degree-of-freedom motion of the upper plate[40]. Given this configuration, we developed a simplified dynamics model under the assumption of a fixed rotational center for the upper plate. Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation. This necessitates adjusting the applied actuating voltages across the three muscle strings to drive the upper plate from its initial equilibrium state to predefined target angles.
The reference frame $O_{\mathrm{xyz}}$ is fixed on the moving upper plate, with the $\mathrm{z}$-axis perpendicular to the upper plate and pointing outward. $\theta,\phi,$ and $\psi$ represent the pitch, roll, and yaw angles, respectively. Similar to the robotic eye, the system state variables are defined as
$\bm{x}=[\theta,\dot{\theta},\phi,\dot{\phi},\Delta T_{1},\Delta T_{2},\Delta T_{3}]^{\top},$ (6)
where $\dot{\theta}$ and $\dot{\phi}$ represent the corresponding angular velocities, and $\Delta T_{i}$ is the temperature difference of muscle $m_{i}$ with respect to the ambient environment, with $i\in\{1,2,3\}$. The system output vector is represented as $\bm{y}=[\theta,\dot{\theta},\phi,\dot{\phi}]^{\top}$. Likewise, the muscle’s temperature is not used as an input for the DRL agent, and thus, the state vector is defined as
$s_{t}=[\theta_{t},\dot{\theta}_{t},\phi_{t},\dot{\phi}_{t},\theta^{*},\phi^{*}]^{\top},$ (7)
where $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate. Additionally, the input action vector $a_{t}$ of the system is defined as:
[…]
The output and control weight matrices of the reward function are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. When the tracking error is sufficiently small, a bonus reward is granted, i.e.,
V Learning Control Framework